Fermi National Laboratory


Accelerator Update for January 28, 2005 - January 31, 2005

Friday January 28
The day shift began with the Tevatron (TeV) in shot setup, with MiniBooNE taking beam, and with the Meson experiment MCenter off due to shot setup.

At 8:44 AM, a TeV sector E sextupole (C:S3E1A) tripped off on a ground fault indication. Operators had its regulator changed by about thirty minutes later.

Operations put the TeV through a dry squeeze at 8:46 AM.

Operations established store 3950 at 11:02 AM with an initial luminosity of 93.07E30. The Antiproton Source (Pbar) resumed stacking.

Operations established beam to Switchyard (SY) and the Meson experiments around 12:44 PM. The delay was due to a timeline problem.

Operations established beam to the Meson MTest experiment at 5:12 PM

Operations put the Recycler into shot setup to stash antiprotons at 9:31 PM and ended at 11:13 PM. Pbar resumed stacking and at 11:41 PM and Operations resumed sending beam to the Meson experiments.

Saturday January 29
The midnight shift began with Operations monitoring store 3950 and the stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

MTest experimenters said that they were through using the beam at 2:29 AM.

At 7:42 AM, Operations reported that a Main Injector (MI) RF station (MIRF18) wasn't putting out the proper RF. Operators contacted an expert who said he'd come it.

MTest began taking beam at 7:51 AM.

An RF expert reported that MIRF18 was back up and running at 9:53 AM. He had found and repaired a faulty cable.

A Meson expert came in at 11:39 AM and worked on tuning SY and the beam lines until 2:40 PM.

Operations terminated store 3950 at 4:23 PM.

Operations put the TeV into shot setup at 5:45 PM.

Operations established store 3952 at 8:10 PM with an initial luminosity of 105.17E30. This set a NEW RECORD LUMINOSITY. Pbar resumed stacking.

Operations resumed sending beam to SY around 8:35 PM.

At 8:37 PM, a Booster RF station (BRF17) tripped off for the first of six times in about an hour. Operators called an expert who suggested turning down the modulator voltage, which helped.

Sunday January 30
The midnight shift began with operations monitoring store 3952, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

At 12:30 AM, the Booster 400 MeV chopper (B:CHOP) tripped off due to a temperature problem. Operators couldn't get the chopper to reset. An expert was called in to look at it. This halted stacking and all downstream beam.

Operations reported that a Booster expert had the chopper back on by 2:10 AM. He improved the airflow.

Recycler stashed antiprotons around 3:30 AM.

Booster experts began tuning the machine at 2:19 PM.

A Pbar expert began a stacking study at 2:58 PM.

At 4:47 PM, the I- Source dome showed a bad status. Operators investigated and found that there was no power to the dome; the brushes for the motor generator had failed. Operators called in Linac experts to help.

Operators and a Linac expert began bringing up the H- Source at 6:16 PM.

Pbar resumed stacking at 6:33 PM with MiniBooNE and the Meson experiments getting beam a few minutes later.

The Linac expert brought the I- Source on line at 6:55 PM.

Monday January 31
The midnight shift began with Operations monitoring store 3952 and the stash, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

At 1:44 AM, the Linac Buncher tripped for the first of two times in eight minutes. It reset okay each time.

At 4:41 AM, Operations stopped stacking to cool the core to prepare for the next shot, but they started having problems with the Pbar VSA. Operators reverted to stacking mode while an expert investigated the VSA set up.

Operation reported at 5:57 AM that they had the VSA back, but were still manually cooling the large emittances.

Operations terminated store 3952 at 5:09 AM and put the TeV into shot setup.

At 5:14 AM, Operations reported that the vacuum in SY enclosure C was bad. Operations turned off the beam. Later a vacuum expert discovered that the problem was due to another vacuum system (G2), which they isolated from the SY system.

Operations reported at 6:42 AM that Pbar was ready to continue with shot setup.

The Plan for Monday
The plan for today is to establish a new store and then continue with stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/31/2005   email Fermilab